simplify: sq rt (-49) +2i^2-10i^7
Answer by edjones(8007)

Remember:
i=i^5=i^9=sqrt(-1)
i^2=i^6=i^10=-1
i^3=i^7=i^11=-i
i^4=i^8=I^12=1
.
sq rt (-49) +2i^2-10i^7
=7i-2+10i
=-2+17i
.
Ed